OpenStudy (geekfromthefutur):

Kylie sells 2 concert tickets every minute. She has already sold 8 tickets. If f(t) represents the total number of tickets that Kylie will sell in t minutes, which of the following functions represents the relationship between f(t) and t? f(t) = 8t - 2 f(t) = 2t - 8 f(t) = 2t + 8 f(t) = 8t + 2

OpenStudy (yanasidlinskiy):

She sells 2 tickets every minute: 2t and and she has already sold 8 tickets: 8 The answer is C: f(t) = 2t + 8.

OpenStudy (geekfromthefutur):

Thx once again lol

OpenStudy (geekfromthefutur):

The equation of a line is shown below. y = - fraction reading 1 over 3 x - 24 What is the equation of a line which is perpendicular to this line and passes through (1, 27) y = 3x - 24 y = 7x - 24 y = 7x + 24 y = 3x + 24